# The Fraction of Clumpy Galaxies in JWST Surveys over $2<z<12$

### v1

@@ 2.2 Photometric and Spectroscopic Redshifts
Figure 1: Comparison of 5,780 spectroscopic and photometric redshifts in the JADES fields. The former are compiled from various sources (see Section 2.2) and the latter are measured using EAZY (Brammer et al., 2008). The bias, normalized median absolute deviation (NMAD) scatter, and outlier fraction are measured according to Dahlen et al. (2013).
The accuracy of the photometric redshifts measured by the JADES team is assessed by comparing them with spectroscopic redshifts. We find 2,774 galaxies with spectroscopic redshifts in GOODS-N and 3,006 in GOODS-S. Their redshifts span $0<z<13.35$. The accuracies of the photometric redshifts are quantified using the bias, normalized median absolute deviation (NMAD) scatter, and outlier fraction according to Dahlen et al. (2013). The bias is given by
@@
Figure 2: The relation between stellar mass and SFR, also known as the star-formation main sequence (SFMS), for the full sample (Section 3.1) over $2<z<12$. These quantities are measured from fits to the combined HST and JWST SEDs using bagpipes (Carnall et al., 2018, see Section 2.3). Measurements for individual galaxies are shown as gray dots in various redshift bins, a different one of which is in each column. Completeness limits in stellar mass are indicated with vertical dashed lines. There is broad agreement between the SFMS presented here and that measured in other studies. The SFMS measured by Popesso et al. (2023) over $2<z<6$ from a compilation of studies is shown as a blue solid line in the leftmost four columns. That measured by Salmon et al. (2015) over $3.5<z<6.5$ is plotted in the second through fourth columns. The SFMS from Cole et al. (2025), measured over $4.5<z<12$, is shown as a green dashed line. The SFMS measurements by Ciesla et al. (2024) over $6<z<12$ are shown as magenta dash-dotted lines.
2.3 Stellar Masses and Star-Formation Rates
Stellar masses and star-formation rates for all galaxies in the full sample (Section 3) are measured by fitting their SEDs. The SFRs used in this work are averaged over the last 100 Myr. Where available, fits were performed to the HST/Advanced Camera for Surveys (ACS) bandpasses F435W, F606W, and F775W combined with all the NIRCam bandpasses listed in Section 2.1. The number of filters used in each fit ranges from five to 17, with a median of 13. The Bayesian SED-fitting tool bagpipes (Carnall et al., 2018) is used. It fits SEDs using the MultiNest Markov Chain Monte Carlo algorithm (Feroz et al., 2009) and returns marginalized posterior probability distribution functions for each parameter of interest. Stellar masses and SFRs are measured by taking the median stellar mass and SFR from their respective posteriors.
@@ 2.3 Stellar Masses and Star-Formation Rates
Priors for other physical parameters are as follows. A uniform prior on the log of the stellar mass $\log\left(M_{\star}/M_{\odot}\right)$ is assumed over the range of 1 to 13. We adopt the flexible attenuation curve model by Salim et al. (2018). It modifies the shape of the attenuation curve measured by Calzetti et al. (2000) by multiplying it by a power law that varies with wavelength. For the slope of the power law, $\delta$, a Gaussian prior is set with $\mu=0$ and $\sigma=0.1$ over the range [-0.3, 0.3]. The Calzetti et al. (2000) curve is obtained by setting $\delta=0$. The strength of the dust absorption feature at rest-frame 2,175 Å is also modeled by Salim et al. (2018). We set a uniform prior on it over the range [0, 5]. Another prior is assumed for the differential attenuation between young ($\leq 10$ Myr) and old stars (see, e.g., Calzetti et al., 2000; Charlot & Fall, 2000). The former are generally dustier than the latter due to additional attenuation from birth clouds. A uniform prior on $\eta$, the ratio of $A_{V}$ for young stars to that for old stars, is assumed over the range [1, 5]. SED fits include nebular emission, the strength of which is determined with the log of the ionization parameter, $\log U$, and the nebular metallicity. A uniform prior is set on the former over the range [4,2-4,-2]. The latter is assumed to be the same as that of the stars.
Last, a double power-law star-formation history is assumed. It is given by
@@
Stellar population synthesis templates from version 2.2.1 of the Binary Population and Spectral Synthesis (BPASS; Eldridge et al., 2017; Stanway & Eldridge, 2018) models are used. The initial mass function is set to the default ‘135_300’ model listed in Table 2 in Stanway & Eldridge (2018). The cloudy (Ferland et al., 2017) photoionization code is used to model nebular emission. All HST and JWST filter transmission curves, which are used in the SED fits, are taken from the SVO Filter Profile Service (Rodrigo et al., 2012; Rodrigo & Solano, 2020).
@@
2.4 Stellar Mass Completeness Limits
@@
The completness limits are estimated as follows. First, a limiting stellar mass, $M_{\rm{lim}}$, is computed for all galaxies in the full sample (Section 3). It is defined to be
@@
where $\log\left(M_{\star}\right)$ is the stellar mass in solar units of a given galaxy, $m$ is its apparent magnitude, and $m_{\rm{lim}}$ is the limiting magnitude of the survey. In this work, the magnitude in the F [^4] W band is used to calculate $M_{\rm{lim}}$. This band is used as it is the longest NIRCam filter available in both GOODS-N and GOODS-S and thus best traces the stellar mass. After finding $M_{\rm{lim}}$ for all galaxies, they are divided into redshift bins (see Section 3.1). In each bin, the faintest 20 percent of galaxies are identified. Then, the 90th percentile of $M_{\rm{lim}}$ for these is found. We take $m_{\rm{lim}}$ to be 29.8, which is the average $5\sigma$ point-source depth across the GOODS-N and GOODS-S fields, as measured by Puskás et al. (2025).
@@
3 Sample Selection
3.1 The Full Sample
We start with all sources in the JADES GOODS-N and JADES GOODS-S photometric catalogs. There are 179,709 sources in both fields: 85,709 in GOODS-N and 94,000 in GOODS-S.
All sources that are brighter than or equal to 29th magnitude in the reddest broadband filter, F [^4] W, are selected. This is about 0.5 mag brighter than the limiting magnitudes for all NIRCam filters in JADES (Table 2 in Rieke et al., 2023b). This cut removes 82,606 objects with 97,103 remaining. Next, we remove point sources using the FLAG_ST flag. This cut removes 531 sources, leaving 96,572 galaxies. All sources that are contaminated by diffraction spikes from bright stars are removed with the FLAG_BS flag. Those that are contaminated by bright neighbors are removed using the FLAG_BN flag. The former cut removes 1,504 galaxies and the latter removes 23,672, with 71,396 left.
@@ 3.1 The Full Sample
Cuts are then applied on the morphologies of galaxies. These are done to exclude compact and edge-on galaxies and have often been applied in pre-JWST studies of clumpy galaxies (e.g., Guo et al., 2012, 2015; Shibuya et al., 2016). Galaxies with circular half-light radii of at least 0.″1 in the F [^4] W bandpass are selected. This removes 2,848 galaxies with 16,345 remaining. Galaxies with axis ratios of $b/a\geq 0.5$ are chosen. This excludes 933 more.
The full sample contains 15,412 galaxies: 7,167 in GOODS-N and 8,245 in GOODS-S. Galaxies in the final sample have a median S/N of about 4.0 in F090W and 13.2 in F277W. For each galaxy in the full sample, its stellar mass and SFR are measured by fitting its combined HST+JWST SED (Section 2.3). About 8% (1,195/15,412) of the full sample has spectroscopic redshifts. The distribution of the full sample in terms of redshift and stellar mass is shown in Figure 4. Individual galaxies are represented with small gray dots. The stellar mass completeness limit for each redshift bin is indicated with a light gray shaded region. Galaxies with spectroscopic redshifts are indicated with empty red squares.
3.2 The Final Sample
Galaxies with stellar masses $\geq 10^{8}~M_{\odot}$ are selected. This is based on our completeness limits, details which are provided in Section 2.4. In the highest redshift bin, $9<z<12$, this cut in mass is about 80% complete. However, we choose to apply it across all redshift bins for consistency. Applying the mass cut removes 6,074 galaxies and leaves 9,338.
Figure 4: The stellar masses as a function of redshift are shown for the full sample (Section 3). Individual galaxies with only photometric redshifts are plotted as small gray dots and those with spectroscopic redshifts are indicated with red squares. Top panel: the distribution of redshifts. Right panel: stellar mass distribution. In the top and right panels, the distributions for galaxies with spectroscopic redshifts are denoted with red dash-dot histograms. Clumps are identified in galaxies with stellar masses of at least $10^{8}~M_{\odot}$, which is indicated with the horizontal dashed line. It is higher than or close to the 90% stellar mass completeness limit of the sample at all redshifts, which are shown as light gray shaded regions and capped with dotted lines.
@@ 3.2 The Final Sample
Figure 5: The completeness of our clump detection algorithm in redshift bins (columns) and stellar mass bins of their host galaxies (thickness of the lines) is shown here. Fake clumps are added to rest-frame NUV images of real galaxies. More than 1,500 galaxies are randomly selected over redshifts $2<z<12$ and stellar masses $8<\log\left(M_{\star}/M_{\odot}\right)<11$, regardless whether they have clumps. The completeness is defined as the fraction of fake clumps recovered relative to the number inserted into each galaxy image. The relation between the ratio of fake clump luminosity to host galaxy luminosity and completeness is used to determine minimum fractional luminosities used to find clumps. A completeness of 50% is denoted with the horizontal dashed lines. Minimum fractional luminosities that meet this completeness are indicated with vertical dash-dot lines. Those for more massive galaxies are denoted with thicker lines. Completeness curves are shown only for ranges in mass that include at least 10 galaxies. Faint clumps are more likely to be detected in higher-mass galaxies at lower redshifts. For galaxies over $9<\log\left(M_{\star}/M_{\odot}\right)<10$, clumps are identified with 50% completeness at 2% of the host luminosity at $z\sim 2.75$ and at 7% of the host luminosity at $z\sim 7.75$. At $z\sim 2.75$ and $\log\left(M_{\star}/M_{\odot}\right)\geq 9$, clumps can be identified at nearly four times fainter luminosities with JWST than with HST. The error on the completeness is calculated using Bayesian binomial confidence intervals and shown as the light shading around each line.

### v2

@@ II.2 Photometric and Spectroscopic Redshifts
Figure 1: Comparison of 5,780 spectroscopic and photometric redshifts in the JADES fields. The former are compiled from various sources (see Section II.2) and the latter are measured using EAZY (Brammer et al., 2008). The bias, normalized median absolute deviation (NMAD) scatter, and outlier fraction are measured according to Dahlen et al. (2013).
The accuracy of the photometric redshifts measured by the JADES team is assessed by comparing them with spectroscopic redshifts. We find 2,774 (3006) with spectroscopic redshifts in GOODS-N (GOODS-S). Their redshifts span $0<z<13.35$. The accuracies of the photometric redshifts are quantified using the bias, normalized median absolute deviation (NMAD) scatter, and outlier fraction according to Dahlen et al. (2013). The bias is given by
@@
Figure 2: The relation between stellar mass and SFR, also known as the star-formation main sequence (SFMS), for the sample of galaxies (Section III) over $2<z<12$. These quantities are measured from fits to the combined HST and JWST SEDs using bagpipes (Carnall et al., 2018, see Section II.3). Measurements for individual galaxies are shown as gray dots in various redshift bins, a different one of which is in each panel. Completeness limits in stellar mass in each redshift bin are indicated with vertical dashed lines. There is broad agreement between the SFMS presented here and that measured in other studies. The SFMS measured by Popesso et al. (2023) over $2<z<6$ from a compilation of studies is shown as a blue solid line in the leftmost four panels. That measured by Salmon et al. (2015) over $3.5<z<6.5$ is plotted in the second through fourth columns. The SFMS from Cole et al. (2025), measured over $4.5<z<12$, is shown as a green dashed line. The SFMS measurements by Ciesla et al. (2024) over $6<z<12$ are shown as magenta dash-dotted lines.
II.3 Stellar Masses and Star-Formation Rates
Stellar masses and star-formation rates for all galaxies in the full sample (Section III) are measured by fitting their SEDs. The SFRs used in this work are taken to be the average values inferred from the most recent 100 Myr in the fit star-formation histories (SFHs). Where available, fits were performed to the HST/Advanced Camera for Surveys (ACS) bandpasses F435W, F606W, and F775W combined with all the NIRCam bandpasses listed in Section II.1. The number of filters used in each fit ranges from five to 17, with a median of 13. The Bayesian SED-fitting tool bagpipes (Carnall et al., 2018) is used. It fits SEDs using the MultiNest Markov Chain Monte Carlo algorithm (Feroz et al., 2009) and returns marginalized posterior probability distribution functions for each parameter of interest. Stellar masses and SFRs are measured by taking the median stellar mass and SFR from their respective posteriors.
@@ II.3 Stellar Masses and Star-Formation Rates
Priors for other physical parameters are as follows. A uniform prior on the log of the stellar mass $\log\left(M_{\star}/M_{\odot}\right)$ is assumed over the range of 1 to 13. We adopt the flexible attenuation curve model by Salim et al. (2018). It modifies the shape of the attenuation curve measured by Calzetti et al. (2000) by multiplying it by a power law that varies with wavelength. For the slope of the power law, $\delta$, a Gaussian prior is set with $\mu=0$ and $\sigma=0.1$ over the range [-0.3, 0.3]. The Calzetti et al. (2000) curve is obtained by setting $\delta=0$. The strength of the dust absorption feature at rest-frame 2,175 Å is also modeled by Salim et al. (2018). We set a uniform prior on it over the range [0, 5]. Another prior is assumed for the differential attenuation between young ($\leq 10$ Myr) and old stars (see, e.g., Calzetti et al., 2000; Charlot and Fall, 2000). The former are generally dustier than the latter due to additional attenuation from birth clouds. A uniform prior on $\eta$, the ratio of $A_{V}$ for young stars to that for old stars, is assumed over the range [1, 5]. SED fits include nebular emission, the strength of which is determined with the log of the ionization parameter, $\log U$, and the nebular metallicity. A uniform prior is set on the former over the range [4,2-4,-2]. The latter is assumed to be the same as that of the stars.
Last, a double power-law SFH is assumed, given by
@@
Stellar population synthesis templates from version 2.2.1 of the Binary Population and Spectral Synthesis (BPASS; Eldridge et al., 2017; Stanway and Eldridge, 2018) models are used. The initial mass function is set to the default ‘135_300’ model listed in Table 2 in Stanway and Eldridge (2018). The cloudy (Ferland et al., 2017) photoionization code is used to model nebular emission. All HST and JWST filter transmission curves, which are used in the SED fits, are taken from the SVO Filter Profile Service (Rodrigo et al., 2012; Rodrigo and Solano, 2020).
@@
II.4 Stellar Mass Completeness Limits
@@
The completness limits are estimated as follows. First, a limiting stellar mass, $M_{\rm{lim}}$, is computed for all galaxies in the full sample (Section III). It is defined to be
@@
where $\log\left(M_{\star}\right)$ is the stellar mass in solar units of a given galaxy, $m$ is its apparent magnitude, and $m_{\rm{lim}}$ is the limiting magnitude of the survey. In this work, the magnitude in the F [^4] W band is used to calculate $M_{\rm{lim}}$. This band is used as it is the longest-wavelength NIRCam filter available in both GOODS-N and GOODS-S and thus best traces the stellar mass. After finding $M_{\rm{lim}}$ for all galaxies, we divide them in redshift bins (Section III). In each bin, the faintest 20 percent of galaxies are identified. Then, the 90th percentile of $M_{\rm{lim}}$ for these is found. We take $m_{\rm{lim}}$ to be 29.8, which is the average $5\sigma$ point-source depth across the GOODS-N and GOODS-S fields, as measured by Puskás et al. (2025).
@@
III Selection of the Sample of Galaxies
We start with all sources in the JADES GOODS-N and GOODS-S photometric catalogs. There are 179,709 sources in both fields: 85,709 in GOODS-N and 94,000 in GOODS-S.
All sources that are brighter than or equal to 29th magnitude in the reddest broadband filter, F [^4] W, are selected. This is about 0.5 mag brighter than the limiting magnitudes for all NIRCam filters in JADES (Table 2 in Rieke et al., 2023a). This cut removes 82,606 objects with 97,103 remaining. Next, we remove point sources using the FLAG_ST flag. This cut removes 531 sources, leaving 96,572 galaxies. All sources that are contaminated by diffraction spikes from bright stars are removed with the FLAG_BS flag. Those that are contaminated by bright neighbors are removed using the FLAG_BN flag. The former cut removes 1,504 galaxies and the latter removes 23,672, with 71,396 left.
@@ III Selection of the Sample of Galaxies
Cuts are then applied to exclude compact and edge-on galaxies, following pre-JWST studies of clumpy galaxies (e.g., Guo et al., 2012, 2015; Shibuya et al., 2016; Sattari et al., 2023). Galaxies with circular half-light radii of at least 0.″1 in the F [^4] W bandpass are selected. This minimum size is slightly larger than the half-width at half maximum of the F [^4] W point spread function (PSF, Zhuang and Shen 2024) and is imposed in order to select galaxies that are resolved. This removes 2,848 galaxies with 16,345 remaining. Galaxies with axis ratios of $b/a\geq 0.5$ are chosen. This excludes 933 more galaxies.
The full sample contains 15,412 galaxies: 7,167 in GOODS-N and 8,245 in GOODS-S. Galaxies in the full sample have a median S/N of about 4.0 in F090W and 13.2 in F277W. For each galaxy in the sample, its stellar mass and SFR are measured by fitting its combined HST+JWST SED (Section II.3). About 8% (1,195/15,412) of the full sample has spectroscopic redshifts. The distribution of the full sample in terms of redshift and stellar mass is shown in Figure 4.
To identify clumps within a mass-complete sample, galaxies with stellar masses $\geq 10^{8}~M_{\odot}$ are selected. This is based on our completeness limits (Section II.4). In the highest redshift bin, $9<z<12$, this cut in mass is about 80% complete. However, we choose to apply it across all redshift bins for consistency. Applying the mass cut removes 6,074 galaxies and leaves 9,338.
Figure 4: The stellar masses as a function of redshift are shown for the full sample (Section III). Individual galaxies with only photometric redshifts are plotted as small gray dots and those with spectroscopic redshifts are indicated with red squares. Top panel: the distribution of redshifts. Right panel: stellar mass distribution. In the top and right panels, the distributions for galaxies with spectroscopic redshifts are denoted with red dash-dot histograms. Clumps are identified in galaxies with stellar masses of at least $10^{8}~M_{\odot}$, which is indicated with the horizontal dashed line. It is higher than or close to the 90% stellar mass completeness limit of the sample at all redshifts, which are shown as light gray shaded regions and capped with dotted lines.
@@
Figure 5: Example clumpy galaxies are shown in each redshift (vertical) and stellar mass bin (horizontal), where a complete sample of clumps is detected. For each galaxy, images in the rest-frame NUV are shown to the left and an RGB image is on the right. Unmodified images are shown to the top left; smoothed images on the top right; contrast images on the bottom left, and detected clumps are circled in pink in the bottom right. Clumps that meet the completeness limits listed in Table 1 are denoted with solid lines, otherwise, with dotted lines. The segmentation map of each galaxy is indicated with the red outline.
Figure 6: Surface brightness profiles of detected clumps. In each panel, the surface brightness profiles of individual clumps are shown as thin gray lines. Those of the PSFs of the filters used to detect clumps are shown with dark blue dotted lines. Median profiles are plotted as thick, black, solid lines and running 16th and 84th percentiles are shown as lower and upper dashed lines. Sérsic profiles are fit to the median profiles and shown as red dash-dot lines. The effective radii and indices of the fits are written in red text. All surface brightness profiles are normalized to that of the innermost aperture used to measure them. Panels are arranged in blocks of 3x3 panels that are indicated with bins in redshift. Only panels with at least 5 detected clumps are shown. Within each block, host galaxy stellar mass increases from left to right and brightness of the clump compared with its host increases from top to bottom. The latter quantity is measured by assuming the clumps behave as point sources. In all panels, clumps are more extended than the PSFs. The angular sizes of clumps are nearly constant with redshift, but their physical sizes increase by a factor of $\sim 2$ from $z\sim 7.75$ to $z\sim 2.75$.
IV Selection of Clumpy Galaxies
IV.1 Identification of Clumps
Clumps are identified for all galaxies in the mass-selected, star-forming sample as follows. For each galaxy in the sample, a 6″ by 6″ postage stamp is created in the bandpass corresponding to the rest-NUV. This image is then smoothed with a boxcar filter of width equal to 10 pixels. Our choice of filter width is justified with injection-recovery tests (Section IV.3). The smoothed image is subtracted from the original image to create a contrast image. Objects are then detected in the contrast image. This is done with the photutils library. The $3\sigma$-clipped standard deviation of the background is measured and source detection is performed. Deblending is done with photutils to separate overlapping clumps, with a minimum contrast of 0.001 and 32 thresholds. Objects are selected to be candidate clumps if they span at least five contiguous pixels, each of which must be at least two standard deviations brighter than the background-subtracted image, and if the total S/N of the object is at least 3. The minimum physical area of a candidate clump is 0.33 kpc2 at $z=2$ and 0.06 kpc2 at $z=12$. To be selected as clumps, candidates must lie at least 0.″1 from the center and within the segmentation map created by the JADES team. This corresponds to a distance of 0.86 kpc at $z=2$ and 0.37 kpc at $z=12$. The former criterion is used to avoid confusion with a bulge or active galactic nucleus and the latter ensures that all detected clumps lie within a single galaxy and not within its neighbors projected on the sky. Examples of the detection technique and identified clumps are shown in Figure 5. The technique and selection criteria above broadly follow many pre-JWST studies (Murata et al., 2014; Guo et al., 2015; Shibuya et al., 2016; Soto et al., 2017; Martin et al., 2023; Sattari et al., 2023).
IV.2 Radial Surface Brightness Profiles and Aperture Photometry of Clumps
In the HST data, clumps were largely found to be point sources (Guo et al., 2015; Shibuya et al., 2016; Sattari et al., 2023). Given the higher spatial resolution of JWST, this assumption may no longer hold. The surface brightness profiles of clumps are shown in Figure 6, divided into bins of redshift, host galaxy stellar mass, and UV brightness of the clumps relative to their host galaxies. Profiles were measured with concentric circular apertures centered on the centroids determined by our clump detection routine.
Clumps detected in this work have spatially extended surface brightness profiles, which are inconsistent with point sources. Empirical PSFs are measured and described in Appendix A. Figure 6 shows that in all bins, the median profiles are more extended than those of the PSFs.
To measure the flux of each clump, a typical surface brightness profile shape is first assumed. Then, the flux is directly measured within a small circular aperture. Background flux from the host galaxy is removed and an aperture correction is performed. This follows previous works (e.g., Guo et al., 2015; Shibuya et al., 2016; Sattari et al., 2023), except for our treatment of the clumps as extended sources.
A Sérsic function is assumed for clump profiles, which is then convolved with the PSF and fit to the median observed profiles (Figure 6). The effective radii and Sérsic indices of the fits are estimated. Relative errors on the effective radii are generally small, of only a few percent. When fitting the median profiles, the background light from the host is not subtracted, but is included in the fit. We assume the light profile over 6106-10 pixels in radius is dominated by the background light, as most profiles flatten out at these radii (see also Guo et al. 2015).
There is little variation in the median clump profile with redshift, host galaxy mass, or clump brightness. Over all bins shown in Figure 6, the median effective radius and Sérsic index from the fits are $R_{\rm{eff}}=$0.″046 and $n=1.04$. The 16–84th percentile range in $R_{\rm{eff}}$ is 0.″044 to 0.″049 and that for $n$ is 0.94 to 1.10. There is a weak evolution in physical size with redshift, such that the median effective radius of clumps grows from 0.23 kpc (0.″045) at $z\sim 7.75$ to 0.39 kpc (0.″048) at $z\sim 2.75$. This is a factor of a few larger than the median size of clumps within lensed galaxies, which was found to be 0.093 kpc over $3.5<z<5.5$ (Claeyssens et al., 2025). The larger sizes reported in this work may be due to our selection of unlensed galaxies. Individual clumps that may only be detectable in lensed galaxies could aggregate into larger ones in unlensed systems.
Aperture photometry of the clumps is performed with photutils as follows. First, the flux of a given clump is measured with a circular aperture of radius 4 pixels. Beyond this radius, the typical clump light profile is dominated by host galaxy light (see Figure 6). The background light is subtracted by measuring the flux in circular annuli over the radii 6-10 pixels. It is then divided by the area of the annulus to yield a mean surface brightness. This is multiplied by the area of the inner aperture with radius of 4 pixels to yield a background flux, which is then subtracted from the original flux measured. In each step, all other clumps are masked with circular masks of radius 4 pixels. The sizes of the apertures, annuli, and masks used here are also the same as those used in Guo et al. (2015).
Last, aperture corrections are applied to background-subtracted fluxes of clumps. These are calculated by assuming clump light profiles broadly follow the median Sérsic profile $(R_{\rm{eff}}=$0.″046, $n=1.04)$. This profile is then convolved with a PSF. The total flux of the convolved profile and that within a circular aperture of radius equal to 4 pixels are measured. The aperture correction is taken to be the ratio of the total flux to that within the aperture. Clump aperture fluxes are multiplied by factors of 1.511, 1.473, 1.477, 1.514, and 1.784 in GOODS-N and 1.533, 1.492, 1.509, 1.530, and 1.738 in GOODS-S for the F090W, F115W, F150W, F200W, and F277W bandpasses, respectively. These are determined using the empirical PSFs (Appendix A).
@@ IV.2 Radial Surface Brightness Profiles and Aperture Photometry of Clumps
Figure 7: Clump completeness functions in redshift (vertical) and stellar mass (horizontal) bins. The smoothing length used to detect clumps are also shown. Completeness is measured by injecting fake clumps to rest-frame NUV images of real galaxies and finding the fraction of fake clumps recovered relative to the number inserted into each galaxy image. The completeness is therefore measured as the ratio of fake clump luminosity to host galaxy luminosity above which >>50% of the fake clumps are detected. A completeness of 50% is denoted with the horizontal dashed lines. Minimum fractional luminosities that meet this completeness are indicated with vertical dash-dot lines. Those for more massive galaxies are denoted with thicker lines. Completeness curves are shown only for ranges in mass that include at least 10 galaxies. The error on the completeness is shown as the light shading around each line.
IV.3 Completeness of Clump Identification
To ensure that a complete sample of clumps is identified, we measure how faint clumps can be in order for half of them to be detected (Guo et al., 2015; Shibuya et al., 2016; Sattari et al., 2023). This is done by adding fake clumps to images of randomly selected real galaxies, regardless whether they have clumps or not, and running the clump identification method described above.
Fake clumps are modeled as Sérsic profiles with the same effective radius and index as the median profile of the clumps. They are convolved with the empirical PSFs measured in this work. The PSF used to convolve the profile varies with redshift as described in Section III. Positions of the fake clumps are randomly selected to lie within the segmentation map of a given galaxy. Their brightnesses vary randomly between 0.1% and 75% of that of the host galaxy. In order to have sufficient statistics to determine the completeness, 100 galaxies are randomly selected in each of three stellar mass bins ($8<\log\left(M_{\star}/M_{\odot}\right)<9,9<\log\left(M_{\star}/M_{\odot}\right)<10$, and $10<\log\left(M_{\star}/M_{\odot}\right)<11$) for each redshift bin. If fewer than 100 galaxies exist within a given bin in mass and redshift, all galaxies within that bin are selected. In total, 1,541 galaxies are randomly selected. A fake clump is added to the rest-frame NUV image of each galaxy. The clump identification algorithm is then run on the modified image to determine whether the fake clump is recovered. For each randomly selected galaxy, this procedure is repeated 50 times. The position and brightness of the fake clumps are randomly varied in each iteration. Uncertainties on the completeness are computed in a Bayesian fashion and assume a noninformative Jeffreys prior on the binomial distribution (Cameron, 2011).
Since our clump detection procedure depends on the choice the smoothing length of the boxcar filter, we perform the above procedure for three choices of smoothing length: 5, 10, and 20 pixels (0.″15, 0.″30, and 0.″60, respectively). If the smoothing length is too small, only the brightest clumps will be detected. If it is too large, spurious clumps may be detected, as noted by Guo et al. (2015). In this work, the optimal length is found to be 10 pixels (see below). It yields maximal completeness while being intermediate in length, thereby minimizing the chance of detecting spurious clumps.
The clump completeness functions are shown in Figure 7. In each panel, the fraction of clumps detected by the clump identification algorithm is shown as a function of the fake clump luminosity. This is further shown as a function of stellar mass (lines of various thicknesses) and redshift (columns). In each row, completeness curves are shown for a certain smoothing length that is used to detect clumps. Clumps are more likely to be detected in higher-mass systems and lower redshifts. Completeness limits are generally the highest for the shortest smoothing length of 5 pixels. They are smallest at the intermediate length of 10 pixels and change little for the longest smoothing length of 20 pixels. The intermediate length is chosen to be optimal. For all lengths considered, no clumps at $z\sim 10.5$ are detected with at least 50% completeness. In the following Sections, we will only focus on the redshift range of $2<z<9$.
Table 1 lists the minimum fractional luminosities as a function of stellar mass and redshift for the optimal smoothing length of 10 pixels. At $z\sim 2.75$ and $\log\left(M_{\star}/M_{\odot}\right)>9$, clump identification is 50% complete at about two times fainter luminosities with JWST (3.5%) than with HST (8%; Shibuya et al. 2016).
A galaxy is defined to be clumpy if it has at least one off-center clump with a rest-frame NUV luminosity that is greater than or equal to the completeness limits listed in Table 1. These depend on stellar mass and redshift.
Figure 8: The fraction of clumpy galaxies, $f_{\rm{clumpy}}$, over 2<z<9{2<z<9}. It is highest at $z\sim 2.75$ and declines by a factor of $\sim 3$ from $z\sim 2.75$ to $z\sim 5.75$. Galaxies are considered clumpy if they have at least one off-center clump. Measurements from this work for galaxies with stellar masses of at least $10^{9}~M_{\odot}$ are shown as black hexagons connected by a solid black line. $f_{\rm{clumpy}}$ measurements for galaxies in GOODS-N (GOODS-S) are shown to the left (right) of the solid black line. The minimum fractional luminosity changes with redshift according to Table 1. We compile measurements from the literature and plot them alongside ours. The results from these studies are largely consistent with each other: $f_{\rm{clumpy}}$ rises gradually from $z\sim 8$ to $z\sim 1.5-2$, when it peaks, and declines at lower redshifts. This is qualitatively consistent with the evolution of the cosmic SFR density, which is shown as a thin blue solid line and has been arbitrarily scaled vertically. The results of this study are consistent with the view that clumpy galaxies became more common at later times. However, as redshift decreases, the $f_{\rm{clumpy}}$ calculated in this work rises more steeply than that measured from the bulk of previous studies. At $z\sim 2.75$ and $z\sim 4.25$, our $f_{\rm{clumpy}}$ measurements are higher than those of most studies by factors of $\sim 1.3-3$.
V Fraction of Clumpy Galaxies and Their Evolution with Redshift
For a galaxy to be clumpy, it must have at least one off-center clump identified in its rest-frame NUV image (Section IV.1). Each clump must have a minimum fractional luminosity relative to its host galaxy. These luminosities are listed in Table 1. The fraction of clumpy galaxies ($f_{\rm{clumpy}}$) is defined to be the number of galaxies with clumps divided by the total number of galaxies in a given bin in redshift and stellar mass. All uncertainties are calculated assuming Bayesian confidence intervals on the binomial distribution, unless otherwise noted.